OpenStudy (anonymous):

a sample of chlorine gas contracts from 550 mL to 376 mL when it is cooled from 63 degrees C to 12 degrees C. if the initial pressure of the chlorine was 725 torr, what is the final pressure of the chlorine?

OpenStudy (aaronq):

you wanna use the ideal gas law: PV=nRT system at certain conditions = system at other conditions \[\frac{ P _{1}V _{1} }{ n _{1}RT _{1} }=\frac{ P _{2}V _{2} }{ n _{2}RT _{2} }\] since n and R are constant you can eliminate them \[\frac{ P _{1}V _{1} }{ T _{1} }=\frac{ P _{2}V _{2} }{ T _{2} }\] substitute in your values and solve for your unknown, algebraically